# Pressure and strain effects on the ab initio electronic structure of La3Ni2O7

### v1

@@ III.1.2 La-5 band lowering and self-doping
Beside this important qualitative modification of the Fermi surface, correlation corrections induce an even more evident and quantitative change in the bandwidth of the 5 band contributed by the two La spacer atoms only (i.e., with no contribution by the La inner atom lying in between the NiO planes, see Fig.Β 13). This La-5 band, which in DFT-PBE is well above the Ni-3 one, is lowered by corrections, as it can be seen at in Fig.Β 1d. At the M point, which is often overlooked in 2D calculations restricted to the plane only, the lowering in is so large (0.7Β eV) that the La-5 crosses the Ni-3 band and becomes the lowest lying conduction band at M. Due to this large correction, the bottom of conduction (BOC) of the La-5 band at M is situated just only 30 meV above the top of valence (TOV) of the Ni-3 band at P (in the following, we will refer to this TOV-BOC distance of 30 meV as the indirect gap). As a consequence, the Fermi level results from a delicate equilibrium between these two incipient bands: the Ni-3 band, which tends to open a pocket of holes at the BZ corners; and the La-5 one, which tends to open a pocket of electrons at M, hereafter labeled . In the latter case, states located on the La atoms of the spacer start to be occupied instead of Ni states. This results in an effective self-doping (SD) by holes of the NiO2 planes. Note that a possible role of rare-earth self-doping was suggested [45] as responsible for the suppression of charge order, thereby possibly favoring superconductivity.
We can expect that any tiny perturbation of the system, e.g.Β doping, vacancies or defects, can drive the system to a qualitatively different electronic structure and Fermi surface, and therefore affect the superconductivity. Note that even computationally, these low-energy features are highly sensitive to the uncertainty of the integration method to calculate the Fermi level, which we discuss in further detail in App.Β C. We mention two recent works which pointed out the shift of the La-5 band due to correlations [15] and highlight the importance of this band on the screening which affects correlations on the Ni- bands [70]. The latter work remarked in particular some coincidences between the lowering trend of the La-5 band with pressure and the shape of the superconducting region. In our calculation this band reaches the Fermi level at lower pressures and crosses the Fermi level already at 30 GPa at the M point, in comparison to 80 Gpa at the point in Verraes et al.Β [70].
@@ III.1.3 Upper-lower Ni-3 energy gap
A feature of the electronic structure that is considered of particular interest in relation to the emergence of superconductivity in La3Ni2O7 is the energy difference between the upper and lower Ni-3 bands. In the literature they are often labeled as bonding- and anti-bonding (as we explain in App.Β H, we find that this labeling is not appropriate). This energy difference is related to the Ni interlayer hopping. According to the scenario proposed by Sakakibara et al.Β [24], a superconductivity gap would open on these two Ni-3 bands even though they can be as far as 1Β eV away from the Fermi level, while the SC gap would be almost zero for the more Ni-3 character bands crossing the Fermi level. In this scenario, a key parameter is precisely this Ni-3 interlayer hopping. As it can be seen in Fig.Β 1d, the two Ni-3 are practically unaffected by corrections, as their band width and separation remain essentially the same as in the DFT calculation. On this point our calculation is in disagreement both with the calculation of Christiansson et al.Β [45], which found a negative downward shift of the upper band of Β eV at the BZ corner, and with the calculation of You et al.Β [15], which in contrast found a positive upward shift of Β eV. Accordingly, our results leave practically unaffected the scenario proposed by Sakakibara et al.Β [24], while it would be boosted or disrupted in the two other cases.
@@
A calculation can also provide the quasiparticle spectral weight which represents a measure of the degree of correlation of the system. In Fig.Β 2a we show the interpolated bandplot of La3Ni2O7 at 29.5Β GPa, with the dots corresponding to the energies sampled on the -mesh. The relative size of the dots indicates the corresponding value of . In Fig.Β 2b we show the general orbital character of the bands (for more detailed projections, see Fig.Β 12). These plots are complemented with Fig.Β 3 which illustrates the behavior of as a function of the energy for the N -point of the Brillouin zone, a point we have chosen for the clear picture it gives. We note that La-5 and 4 bands, which at the level are difficult to distinguish, present the maximum QP weight culminating at . This is already quite remarkable, in particular for the most localized flat bands. Some of the La bands, in particular the previously discussed La-5 which almost achieved the Fermi level, bring these high spectral weights lower in energy than the main La manifold located around 5Β eV. For La-5 states above 5 eV, the QP weight is slightly decreasing (see Fig.Β 3). On the other side of the Fermi level, we also observe for the O-2 manifold below Β eV a reduced with increasing distance from with values ranging from 0.75 to 0.70. This is surprisingly lower than for La-5 and 4 states. However, the most remarkable fact is that the Ni-3 bands, with no difference between and , present the lowest weight of all quasiparticle states. For these states we observe a non-negligible drop of 0.15 when compared in particular with the La-5 which at M is situated at the same Fermi level as the Ni-3. This fact points to a larger weight of the non-coherent part of the spectrum for the Ni-3 states.
@@
We remind that in a Fermi liquid [59] the general trend is a reduction of the quasiparticle weight when going from the Fermi level to lower energies. Indeed, when going to the lowest energy quasiparticle states, the non-coherent part of spectra due to, e.g.Β plasmon losses, should increase, and correspondingly decreases. The same happens in the other higher energies direction, although not monotonically at the beginning. From this, we can conclude that already with the correlations introduced by a non-self-consistent approximation, La3Ni2O7 show a Fermi liquid trend for La-5, La-4 and O-2 quasiparticle states, but not for the Ni-3 ones around the Fermi level. The behaviour of the latter shows a departure from the Fermi liquid trend, but the minimal value of is still large enough not to configure a full breakdown of the Fermi liquid towards a strongly-correlated picture.
@@
As shown in Figs.Β 1a and b, the Fermi surface of La3Ni2O7 exhibits negligible dispersion along the direction, highlighting the pronounced 2D character of its electronic structure. We note that, while this situation is similar to what is found in most of the cuprates, it differs from the infinite-layer nickelates case that presents large dispersion which introduces 3D features in their Fermi surface [71, 72]. Moreover, we note that the dispersion of the Fermi surface of La3Ni2O7 along and is also very weak. In fact, the and Fermi surface sheets can effectively be seen as a superposition of two sets of parallel planes along and , respectively. That is, as the superposition of 1D Fermi surfaces. These effectively 1D Fermi surfaces is an important difference with respect to cuprates. In particular, the emergence of or spin-density-wave ordersβas opposed to the antiferromagnetic order characteristic of the cupratesβcan be related to their nesting properties [34].
@@
In the cuprates, ARPES experiments have revealed the emergence of so-called βFermi arcsβ. Specifically, the quasiparticle spectral weight is found to be locally suppressed at points along the Fermi surface that would coincide with the 1D Fermi surface segments. Unfortunately, to check whether a similar phenomenon takes place in La3Ni2O7, the 666 -sampling we have used in our calculation is not precise enough. Nevertheless, any experimental indication along this direction might provide a clue to understand superconductivity in both nickelates and cuprates.
@@ III.2 Evolution with pressure
In order to gain further insight on the key aspects that might be relevant for superconductivity in La3Ni2O7, we also analyze the evolution of its electronic structure with pressure. We remind that La3Ni2O7 enters into a superconducting phase at an experimentally measured pressure of 14Β GPa without doping [1].
@@
To explore the largest range, we have included in our study a fictitious zero pressure phase of La3Ni2O7 enforced in the tetragonal I4/mmm crystal structure, which in reality at 0Β GPa should be unstable towards the orthorhombic Amam phase with tilted octahedra. Of course, at the present knowledge we cannot exclude that the structural transition occurs at the same critical pressure as the transition to superconductivity, so that it is an important ingredient to explain the superconductivity in this system. In this case, our calculation of the tetragonal phase at 0Β GPa would be probably inadequate to study the normal-superconducting phase transition. But beyond the fact that a calculation is less cumbersome in systems with more symmetries, our choice is motivated by the fact that the comparison of electronic structures is difficult between different crystal structures. Topological variations in the Fermi surface might trigger the superconducting phase transition (and maybe also the structural one): the approach proposed in this section focus on them.
@@
In Fig.Β 4 we compare the electronic structure of La3Ni2O7 at ambient pressure and at 29.5Β GPa (in the same figure we also report the band plot of a biaxial strained La3Ni2O7 which will be discussed in the next section). The most remarkable difference between the two pressures is in the La-5 band at the point M which, at ambient pressure, lies eV above the upper Ni-3 band in a very similar way to what already seen but at the DFT-PBE level in the 29.5 GPa case (see Fig.Β 1d), though in the latter the effect is much larger (0.5 eV). That is, correlations and pressure act in the same direction towards the emergence of the La-5 band at the lowest energies. For a more complete discussion on this point, see App.Β B. At ambient pressure the La-5 band is not present in the low energy region, and therefore the delicate balance between the two incipient self-doping bands is lost.
@@
The electronic structure of La3Ni2O7 at ambient pressure has been measured by an ARPES experiment [56] which we can now compare with our calculation at zero GPa. The hole pocket, which is found in the DFT PBE Fermi surface as due to the lower Ni-3 band, is absent in the ARPES experiment, exactly like in (Fig.Β 1b and Fig.Β 16 in the Appendix). ARPES spectra situate the top of this band at the BZ corners and 50Β meV below the Fermi level. Our calculation at zero GPa (see Fig.Β 4), found the top of this band 50Β meV below the Fermi level at the X -point (corresponding to in Yang et al.Β [56], see Fig.Β 1c for -points correspondence). In contrast to the La-5 band, this lower Ni-3 band is almost unaffected by hydrostatic pressure: in the results at 29.5Β GPa it just only shifts down Β meV, which validates the comparison of the latter case with the ARPES, even if not performed at the same pressure.
@@
In Fig.Β 4b, we report a graph presenting the top-of-valence (TOV) and bottom-of-conduction (BOC) positions of the most relevant bands for different hydrostatic pressures. We report extra calculations done at 14 and at 40Β GPa, covering the range of superconductivity as measured experimentally in Sun et al.Β [1]. The indirect gap is at its minimum around 29.5Β GPa, and the La-5 band is clearly self-doping beyond. From the same figure we can also remark that the same La-5 band crosses the Ni-3 upper band precisely at 14Β GPa, which is precisely the onset of the superconducting region measured by Sun et al.Β [1]. This might be just only a coincidence or have deeper implications. In support to the latter, we cite a very recent work [70] which also found a correlation between the position of the La-5 band with respect to the Fermi level, and the shape of the superconducting region. Their calculation is at the DFT level, so that they missed the earlier coincidence in pressure of the La-5 with the Ni-3 band precisely at the onset of superconductivity at 14Β GPa, but the relevance of the La-5 for superconductivity has been remarked. We can more carefully conclude that both correlations and pressure are fundamental to describe the electronic structure of La3Ni2O7 in its superconducting phase. In the next section we will see that -plane strain plays a similar role to correlations and pressure.
@@ III.3 La3Ni2O7 under epitaxial strain
Recent works [5, 6] have reported superconductivity at Β K in La3Ni2O7 thin films under compressive strain at ambient pressure. Here we present the electronic structure of a bulk La3Ni2O7 with a 1.8% strained Β Γ Β and a relaxed , which we use to simulate the experimental thin film. The corresponding PBE and band plots, PDOS and DOS are shown in Fig.Β 5. In comparison to the previously discussed case at ambient and 29.5Β GPa (see also Fig.Β 4), the electronic structure retains most of its features. The Fermi level crossing is practically the same in the three different cases. In particular, the cuprate-like sheet has exactly the same shape (refer to Fig.Β 1). The lower Ni-3 band is significantly further from in the thin film case ( eV). However, we observe two important differences. First, the hole pocket is already absent at the DFT level, that is the lower Ni-3 band does not cross any longer, an effect that is further enhanced by . Second, the La-5 band at M, which also undergoes a large negative correction of almost 1Β eV, now stays above the Fermi level by eV. In fact, it only reaches the upper Ni-3 band without crossing it. We can see that, with respect to the band positions at M, the epitaxial strained configuration is intermediate between 29.5Β GPa and ambient pressure. Although our biaxial strain simulation implements a non-hydrostatic realization, we can compare it to the hydrostatic pressure case where the La-5 BOC is degenerate with Ni-3 BOC at M. Following this analogy, from Fig.Β 4b, where we have also reported the band positions for the thin film, we can tentatively attribute our epitaxial strain simulation as equivalent to the 14Β GPa hydrostatic pressure case, that is precisely at the superconducting onset. This could be an indication that the relative position and overlap of BOC at M, might play a role as a triggering mechanism of superconductivity.
@@
ARPES spectra have been measured on 1, 2 and 3 unit-cell epitaxial La2.85Pr0.15Ni2O7 films grown on SrLaAlO4 substrates [57]. It is claimed that, in addition to the and sheets, the pocket is also present at the Fermi level. In another more recent ARPES experiment on a La2PrNi2O7 thin film [73], the pocket is absent from the Fermi level and the Ni-3 band is found 70Β meV below the Fermi energy, which compares more favorably with our result of 200Β meV. In any case, the comparison with our electronic structure is difficult because they considered a Pr doped system. From our side, we did not include the neighborhood effect of the substrate which is certainly affecting the electronic structure of such thin films. Our results should then be considered as an ab initio prediction of the electronic structure which could be measured for a La3Ni2O7 thin film of enough large thickness. Or, possibly, for a non-hydrostatic planar strain applied on a bulk sample that can be realized experimentally by applying a tensile strain to the axis, which will induce a biaxial compressive strain of the plane by Poisson law.
@@
We can conclude that our calculation on the epitaxial strained case presents an electronic structure very similar to the bulk at 14Β GPa pressure. Correlation effects are very important on both the lower Ni-3 band, causing the disappearance of the pocket, and on the increased relevance of the La-5 band at low energy.

### v2

@@ III.1.2 La-5 band lowering and self-doping
Beside this important qualitative modification of the Fermi surface, correlation corrections induce an even more evident and quantitative change in the bandwidth of the 5 band contributed by the two La spacer atoms only (i.e., with no contribution by the La inner atom lying in between the NiO planes, see Fig. 13). This La-5 band, which in DFT-PBE is well above the Ni-3 one, is lowered by corrections, as it can be seen at in Fig. 1d. At the M point, which is often overlooked in 2D calculations restricted to the plane only, the lowering in is so large (0.7 eV) that the La-5 crosses the Ni-3 band and becomes the lowest lying conduction band at M. Due to this large correction, the bottom of conduction (BOC) of the La-5 band at M is situated just only 30 meV above the top of valence (TOV) of the Ni-3 band at P (in the following, we will refer to this TOV-BOC distance of 30 meV as the indirect gap). As a consequence, the Fermi level results from a delicate equilibrium between these two incipient bands: the Ni-3 band, which tends to open a pocket of holes at the BZ corners; and the La-5 one, which tends to open a pocket of electrons at M, hereafter labeled . In the latter case, states located on the La atoms of the spacer start to be occupied instead of Ni states. This results in an effective self-doping (SD) by holes of the NiO2 planes. Note that a possible role of rare-earth self-doping was suggested [45] as responsible for the suppression of charge order, thereby possibly favoring superconductivity.
We can expect that any tiny perturbation of the system, e.g. doping, vacancies or defects, can drive the system to a qualitatively different electronic structure and Fermi surface, and therefore affect the superconductivity. Note that even computationally, these low-energy features are highly sensitive to the uncertainty of the integration method to calculate the Fermi level, which we discuss in further detail in App. D. We mention two recent works which pointed out the shift of the La-5 band due to correlations [15] and highlight the importance of this band on the screening which affects correlations on the Ni- bands [70]. The latter work remarked in particular some coincidences between the lowering trend of the La-5 band with pressure and the shape of the superconducting region. In our calculation this band reaches the Fermi level at lower pressures and crosses the Fermi level already at 30 GPa at the M point, in comparison to 80 Gpa at the point in Verraes et al. [70].
@@ III.1.3 Upper-lower Ni-3 energy gap
A feature of the electronic structure that is considered of particular interest in relation to the emergence of superconductivity in La3Ni2O7 is the energy difference between the upper and lower Ni-3 bands. In the literature they are often labeled as bonding- and antibonding (as we explain in App. I, we find that this labeling is not appropriate). This energy difference is related to the Ni interlayer hopping. According to the scenario proposed by Sakakibara et al. [24], a superconductivity gap would open on these two Ni-3 bands even though they can be as far as 1 eV away from the Fermi level, while the SC gap would be almost zero for the more Ni-3 character bands crossing the Fermi level. In this scenario, a key parameter is precisely this Ni-3 interlayer hopping. As it can be seen in Fig. 1d, the two Ni-3 are practically unaffected by corrections, as their band width and separation remain essentially the same as in the DFT calculation. On this point our calculation is in disagreement both with the calculation of Christiansson et al. [45], which found a negative downward shift of the upper band of eV at the BZ corner, and with the calculation of You et al. [15], which in contrast found a positive upward shift of eV. Accordingly, our results leave practically unaffected the scenario proposed by Sakakibara et al. [24], while it would be boosted or disrupted in the two other cases.
@@
A calculation can also provide the quasiparticle spectral weight which represents a measure of the degree of correlation of the system. In Fig. 2a we show the interpolated bandplot of La3Ni2O7 at 29.5 GPa, with the dots corresponding to the energies sampled on the -mesh. The relative size of the dots indicates the corresponding value of . In Fig. 2b we show the general orbital character of the bands (for more detailed projections, see Fig. 15). These plots are complemented with Fig. 3 which illustrates the behavior of as a function of the energy.
@@
We note that La-5 bands, followed by the 4, present the maximum QP weight culminating at . This is already quite remarkable, in particular for the most localized flat bands. Some of the La bands, in particular the previously discussed La-5 which almost achieved the Fermi level, bring these high spectral weights lower in energy than the main La manifold located around 5 eV. For La-5 states above 5 eV, the QP weight is slightly decreasing. On the other side of the Fermi level, we also observe for the O-2 manifold below eV a reduced with increasing distance from with values ranging from 0.75 to 0.70. This is surprisingly lower than for La-5 and 4 states. However, the most remarkable fact is that the Ni-3 bands, with no difference between and , present the lowest weight of all quasiparticle states. For these states we observe a nonnegligible drop of up to 0.15 when compared in particular with the La-5 which at M is situated at the same Fermi level as the Ni-3. This fact points to a larger weight of the noncoherent part of the spectrum for the Ni-3 states.
@@
For comparison, in Fig. 3 we also report the of the jellium model as in the Hedin calculation at and 2 [59], as well as at the La3Ni2O7 nominal average density of at 29.5 GPa ( with a.u., the unit cell volume, and the number of correlated electrons entering into the self-energy from the chosen pseudopotentials). The general trend for a Fermi liquid is a reduction of the quasiparticle weight when going from the Fermi level to lower energies. Indeed, when going to the lowest energy quasiparticle states, the noncoherent part of spectra due to, e.g. plasmon losses, should increase, and correspondingly decreases. The same happens in the other higher energies direction, although not monotonically at the beginning (for and 1.29, this decrease of occurs outside the plot range). From this, we can conclude that already with the correlations introduced by a non self-consistent approximation, La3Ni2O7 show a Fermi liquid trend for all quasiparticle states. We can associate the O-2 and even the La-5 and La-4 quasiparticle weights to the curve, that is to a higher level of correlation with respect to the system nominal density. On their side, the Ni-3 states show a departure from this Fermi liquid trend and so a even higher degree of correlation, although the minimal value of is still large enough not to configure a full breakdown of the Fermi liquid towards a strongly correlated picture.
Of course, this is the picture emerging from a calculation which could be confuted by future experimental measures of the . Should a Fermi liquid breakdown be measured, we will have a failure of the approximation pointing to the need to go beyond it.
@@
As shown in Figs. 1a and b, the Fermi surface of La3Ni2O7 exhibits negligible dispersion along the direction, highlighting the pronounced 2D character of its electronic structure. We note that, while this situation is similar to what is found in most of the cuprates, it differs from the infinite-layer nickelates case that presents large dispersion which introduces 3D features in their Fermi surface [71, 72]. Moreover, we note that the dispersion of the Fermi surface of La3Ni2O7 along and is also very weak. In fact, the and Fermi surface sheets can effectively be seen as a superposition of two sets of parallel planes along and , respectively. That is, as the superposition of 1D Fermi surfaces. These effectively 1D Fermi surfaces is an important difference with respect to cuprates. In particular, the emergence of or spin-density-wave orders—as opposed to the antiferromagnetic order characteristic of the cuprates—can be related to their nesting properties [34].
@@
In the cuprates, ARPES experiments have revealed the emergence of so-called “Fermi arcs”. Specifically, the quasiparticle spectral weight is found to be locally suppressed at points along the Fermi surface that would coincide with the 1D Fermi surface segments. Unfortunately, to check whether a similar phenomenon takes place in La3Ni2O7, the 666 -sampling we have used in our calculation is not precise enough. Nevertheless, any experimental indication along this direction might provide a clue to understand superconductivity in both nickelates and cuprates.
@@ III.2 Evolution with pressure
To gain further insight on the key aspects that might be relevant for superconductivity in La3Ni2O7, we also analyze the evolution of its electronic structure with pressure. We remind that La3Ni2O7 enters into a superconducting phase at an experimentally measured pressure of 14 GPa without doping [1].
@@
To explore the largest range, we have included in our study a fictitious zero pressure phase of La3Ni2O7 enforced in the tetragonal I4/mmm crystal structure, which in reality at 0 GPa should be unstable towards the orthorhombic Amam phase with tilted octahedra. Of course, at the present knowledge we cannot exclude that the structural transition occurs at the same critical pressure as the transition to superconductivity, so that it is an important ingredient to explain the superconductivity in this system. In this case, our calculation of the tetragonal phase at 0 GPa would be probably inadequate to study the normal-superconducting phase transition. But beyond the fact that a calculation is less cumbersome in systems with more symmetries, our choice is motivated by the fact that the comparison of electronic structures is difficult between different crystal structures. Topological variations in the Fermi surface might trigger the superconducting phase transition (and maybe also the structural one): the approach proposed in this section focus on them.
@@
In Fig. 4 we compare the electronic structure of La3Ni2O7 at ambient pressure and at 29.5 GPa (in the same figure we also report the band plot of a biaxial strained La3Ni2O7 which will be discussed in the next section). The most remarkable difference between the two pressures is in the La-5 band at the point M which, at ambient pressure, lies eV above the upper Ni-3 band in a very similar way to what already seen but at the DFT-PBE level in the 29.5 GPa case (see Fig. 1d), though in the latter the effect is much larger (0.5 eV). That is, correlations and pressure act in the same direction towards the emergence of the La-5 band at the lowest energies. For a more complete discussion on this point, see App. B. At ambient pressure the La-5 band is not present in the low energy region, and therefore the delicate balance between the two incipient self-doping bands is lost.
@@
The electronic structure of La3Ni2O7 at ambient pressure has been measured by an ARPES experiment [56] which we can now compare with our calculation at zero GPa. The hole pocket, which is found in the DFT PBE Fermi surface as due to the lower Ni-3 band, is absent in the ARPES experiment, exactly like in (Fig. 1b and Fig. 17 in the Appendix). ARPES spectra situate the top of this band at the BZ corners and 50 meV below the Fermi level. Our calculation at zero GPa (see Fig. 4), found the top of this band 50 meV below the Fermi level at the X -point (corresponding to in Yang et al. [56], see Fig. 1c for -points correspondence). In contrast to the La-5 band, this lower Ni-3 band is almost unaffected by hydrostatic pressure: in the results at 29.5 GPa it just only shifts down meV, which validates the comparison of the latter case with the ARPES, even if not performed at the same pressure.
@@
In Fig. 4b, we report a graph presenting the top-of-valence (TOV) and bottom-of-conduction (BOC) positions of the most relevant bands for different hydrostatic pressures. We report extra calculations done at 14 and at 40 GPa, covering the range of superconductivity as measured experimentally in Sun et al. [1]. The indirect gap is at its minimum around 29.5 GPa, and the La-5 band is clearly self-doping beyond. From the same figure we can also remark that the same La-5 band crosses the Ni-3 upper band precisely at 14 GPa, which is precisely the onset of the superconducting region measured by Sun et al. [1]. This might be just only a coincidence or have deeper implications. (Note that the position of this crossing and the overall picture does not change by performing a calculation in the ambient and low pressure Amam structure which will place the position of this band at even larger energy with respect to its I4/mmm position shown in Fig. 4 at 0 GPa, see App. C). In support to the latter, we cite a very recent work [70] which also found a correlation between the position of the La-5 band with respect to the Fermi level, and the shape of the superconducting region. Their calculation is at the DFT level, so that they missed the earlier coincidence in pressure of the La-5 with the Ni-3 band precisely at the onset of superconductivity at 14 GPa, but the relevance of the La-5 for superconductivity has been remarked. We can more carefully conclude that both correlations and pressure are fundamental to describe the electronic structure of La3Ni2O7 in its superconducting phase. In the next section we will see that -plane strain plays a similar role to correlations and pressure.
@@ III.3 La3Ni2O7 under epitaxial strain
Recent works [5, 6] have reported superconductivity at K in La3Ni2O7 thin films under compressive strain at ambient pressure. Here we present the electronic structure of a bulk La3Ni2O7 with a 1.8% strained Å and a relaxed , which we use to simulate the experimental thin film. The corresponding PBE and band plots, PDOS and DOS are shown in Fig. 5. In comparison to the previously discussed case at ambient and 29.5 GPa (see also Fig. 4), the electronic structure retains most of its features. The Fermi level crossing is practically the same in the three different cases. In particular, the cuprate-like sheet has exactly the same shape (refer to Fig. 1). The lower Ni-3 band is significantly further from in the thin film case ( eV). However, we observe two important differences. First, the hole pocket is already absent at the DFT level, that is the lower Ni-3 band does not cross any longer, an effect that is further enhanced by . Second, the La-5 band at M, which also undergoes a large negative correction of almost 1 eV, now stays above the Fermi level by eV. In fact, it only reaches the upper Ni-3 band without crossing it. We can see that, with respect to the band positions at M, the epitaxial strained configuration is intermediate between 29.5 GPa and ambient pressure. Although our biaxial strain simulation implements a nonhydrostatic realization, we can compare it to the hydrostatic pressure case where the La-5 BOC is degenerate with Ni-3 BOC at M. Following this analogy, from Fig. 4b, where we have also reported the band positions for the thin film, we can tentatively attribute our epitaxial strain simulation as equivalent to the 14 GPa hydrostatic pressure case, that is precisely at the superconducting onset. This could be an indication that the relative position and overlap of BOC at M, might play a role as a triggering mechanism of superconductivity.
@@
ARPES spectra have been measured on 1, 2 and 3 unit-cell epitaxial La2.85Pr0.15Ni2O7 films grown on SrLaAlO4 substrates [57]. It is claimed that, in addition to the and sheets, the pocket is also present at the Fermi level. In another more recent ARPES experiment on a La2PrNi2O7 thin film [73], the pocket is absent from the Fermi level and the Ni-3 band is found 70 meV below the Fermi energy, which compares more favorably with our result of 200 meV. In any case, the comparison with our electronic structure is difficult because they considered a Pr doped system. From our side, we did not include the neighborhood effect of the substrate which is certainly affecting the electronic structure of such thin films. Our results should then be considered as an ab initio prediction of the electronic structure which could be measured for a La3Ni2O7 thin film of enough large thickness. Or, possibly, for a nonhydrostatic planar strain applied on a bulk sample that can be realized experimentally by applying a tensile strain to the axis, which will induce a biaxial compressive strain of the plane by Poisson law.
@@
We can conclude that our calculation on the epitaxial strained case presents an electronic structure very similar to the bulk at 14 GPa pressure. Correlation effects are very important on both the lower Ni-3 band, causing the disappearance of the pocket, and on the increased relevance of the La-5 band at low energy.